A flagpole is secured on opposite sides by two guy wires, each of which is 5 ft. longer than the pole. The distance between the points where the wires are fixed to the ground is equal to the length of one guy wire. How tall is the flagpole (to the nearest inch)?
Answer by Photonjohn(42)

You have an equalateral triangle of x+5 on each side.
Using the pythagorean formula where (x+5) is the hypotenuse,
one side is x the height of the pole, and (x+5)/2, which is the side opposite the 30 degree angle, or 1/2 the hypotenuse,
(x=5)^2 = x^2 + ((x+5)/2)^2
solving for x = 5/(.414) = 12.077' or 144.9"